OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = -1 over 2x + 5. A) y = -2x + 2 B) y= 2x - 14 C) y= 1 over 2x - 8 D) y = -1 over 2x - 4

OpenStudy (mertsj):

\[y=\frac{-1}{2x+5}\]

OpenStudy (mertsj):

That is not a line so there is no line that is parallel to it.

jimthompson5910 (jim_thompson5910):

The slope of y = -1/2x + 5 is -1/2 So the slope of the parallel line is -1/2 Now use y = mx+b to find the equation of the parallel line You'll be plugging in m = -1/2 x = 4 y = -6 \[\Large y = mx+b\] \[\Large -6 = \left(\frac{-1}{2}\right)\left(4\right)+b\] \[\Large -6 = \frac{-4}{2}+b\] \[\Large -6 = -2+b\] \[\Large -6+2 = b\] \[\Large -4 = b\] \[\Large b = -4\] So the equation of the parallel line is \[\Large y = -\frac{1}{2}x-4\]

jimthompson5910 (jim_thompson5910):

So the answer is choice D

OpenStudy (anonymous):

Oh my gosh, thank you!

jimthompson5910 (jim_thompson5910):

You're welcome